# A Lagrangian Approach to the Inhomogeneous Incompressible Euler Equation

### v2

@@ 2.2 Geodesic Equation
Now let $X$ be a geodesic on infinite dimensional Riemannian manifold $(\mathsf{SDiff}(D),\mathsf{G})$. With the above computation, D’Alembert principle implies that $\ddot{X}\in T_{X}^{\perp}\mathsf{SDiff}(D)$, hence there exists a scalar function $\lambda:[0,T]\times D\to\mathbb{R}$ such that for any $x\in D$:
@@
$(\nabla^{*}X)^{-1}(x)\int_{D}\nabla^{*}X(z)\nabla p(X(z))\delta_{x}(z)dz=-\nabla p(t,X).$
Which, in Eulerian coordinates, reads (1).
@@
Differentiating (14) in time and denoting by $u=\dot{X}_{t}\circ X_{t}^{-1}$, we have
@@
The above geodesic equation in Eulerian coordinate reads:
@@ 2.3 Second Fundamental Form and Curvature Tensor
Proposition 2.4
Let $\rho\in\mathsf{Dens}(D)$ and $w\in\mathfrak{X}(D)$, then the following statements hold:
  • (i) There exists a unique $v\in\mathfrak{X}_{\sigma}$ and $p\in C^{\infty}(D)$ such that
    $w=\rho v+\nabla p$ (17)
    moreover, $v=(\mathbf{Id}-\mathbf{Q}_{\rho})(\rho^{-1}w)=:\mathbf{P}_{\rho}(\rho^{-1}w)$
  • (ii) Let $\mathcal{H}_{\rho}:=\{\rho^{-1}\nabla\varphi:\varphi\in C^{\infty}(D)\}$, then we have $\mathsf{Ker}(\mathbf{Id}-\mathbf{Q}_{\rho})=\mathcal{H}_{\rho}$.
  • (iii) $\mathsf{Ran}(\mathbf{P}_{\rho})\perp\mathsf{Ker}(\mathbf{P}_{\rho})$ in the sense of $L_{\rho}^{2}$.
Proof. (i) Given $w\in\mathfrak{X}(D)$, for any given $\rho\in\mathsf{Dens}(D)$ we write
$w=\rho[\rho^{-1}w-\rho^{-1}\nabla\Delta_{\rho}^{-1}\nabla\cdot(\rho^{-1}w)]+\nabla\Delta_{\rho}^{-1}\nabla\cdot(\rho^{-1}w)=\rho\mathbf{P}_{\rho}(\rho^{-1}w)+\rho\mathbf{Q}_{\rho}(\rho^{-1}w)$ (18)
Let $v=\mathbf{P}_{\rho}(\rho^{-1}w)$ and $p=\Delta_{\rho}^{-1}\nabla\cdot(\rho^{-1}w)$ we find the decomposition. Moreover, assume $w=\rho v^{\prime}+\nabla p^{\prime}$ for some $v^{\prime}\in\mathfrak{X}(D)$ and $p^{\prime}\in C^{\infty}(D)$, then $p^{\prime}$ can be recovered from solving the following elliptic problem:
@@
Ellipticity of $\Delta_{\rho}$ implies the uniqueness of solution $p^{\prime}$, hence the uniqueness of $v^{\prime}$.
(ii) Picking any $\varphi\in C^{\infty}(D)$, we compute $\mathbf{P}_{\rho}(\rho^{-1}\nabla\varphi)$ as follows:
$\mathbf{P}_{\rho}(\rho^{-1}\nabla\varphi)=\rho^{-1}\nabla\varphi-\rho^{-1}\nabla\Delta_{\rho}^{-1}\nabla\cdot(\rho^{-1}\nabla\varphi)=0.$
Which guaranties $\mathcal{H}_{\rho}\subset\mathsf{Ker}(\mathbf{P}_{\rho})$. Conversely, assume we have $\mathbf{P}_{\rho}\xi=0$, then
$0=\nabla\times(\rho\mathbf{P}_{\rho}\xi)=\nabla\times(\rho\xi)-\nabla\times(\nabla\Delta_{\rho}^{-1}\nabla\cdot\xi)=\nabla\times(\rho\xi).$
which ensures that $\xi\in\mathcal{H}_{\rho}$ and completes the proof of (ii).
(iii) The orthogonality simply follows from the fact that $\mathsf{Ran}(\mathbf{P}_{\rho})\subset\mathfrak{X}_{\sigma}(D)$, while for generic $v\in\mathfrak{X}_{\sigma}(D)$ we have:
$\langle v,\rho^{-1}\nabla\varphi\rangle_{L_{\rho}^{2}}=\langle v,\nabla\varphi\rangle_{2}=0.$
@@
Theorem 2.5
Then the second fundamental form $\Pi_{X}$ of $\mathsf{SDiff}(D)$ at $X$ is given by
@@
Proof. Formula (19) follows from straightforward computation. By definition:
@@
Remark 2.6
Above $\Pi_{X}$ is symmetric on $T_{X}\mathsf{SDiff}(D)\times T_{X}\mathsf{SDiff}(D)$, since
@@
Here we use the fact that $\nabla\cdot[u,v]=0$ provided $u,v\in\mathfrak{X}_{\sigma}(D)$.
@@
$\mathcal{C}_{\xi,\eta}=\int_{D}\Pi_{X}(\xi,\xi)(z)\cdot\Pi_{X}(\eta,\eta)(z)-\lvert\Pi_{X}(\xi,\eta)(z)\rvert^{2}\rho_{0}(z)dz$
@@
Remark 2.7
The sign of the above curvature tensor fails to detect Lagrangian instability of more general class of steady states due to the following reason: Pressureless steady states of homogeneous Euler are essentially shear flows: Let $X$ be the flow of pressureless steady state $u$, by balance of momentum we have for any initial position $a\in D$:
$\ddot{X}(t,a)=(\partial_{t}u+u\cdot\nabla u)(X_{t}(a))=0.$
which implies that particle trajectories are all straight lines.
2.4 A Gibbs-Appell Formulation
Different from the celebrated Lagrange’s least action principle and Hamilton’s stationary action principle in analytical mechanics, there is a much lesser popular principle/formalism in classic mechanics, known as Gibbs-Appell formulation, which connects to the so-called Gauss’s principle of least curvature and Hertz’s principle of least constraint (see [[Lew96]]). The natural playground for this formulation is a mechanical system with ideal constraints, which is compatible with the nature of incompressible fluids. We refer to [[TG23]] for the study of the Navier-Stokes equation from this point of view. Instead of minimizing Lagrangian action under the constraint of fixed endpoints, the Gibbs-Appell principle minimizes the so-called Appellian functional, which is a quadratic function in acceleration, at each instant.
In the context of IIE, we prove the following:
Theorem 2.8
Define the Appell function
$\mathscr{A}(\ddot{X})=\frac{1}{2}\int_{D}\lvert\ddot{X}(t,a)\rvert\rho_{0}(da).$ (20)
for $X\in C^{2}([0,T];\mathsf{Diff}(D))$. Then minimizer $X_{*}$ of $\mathscr{A}$ at each instant $t\in[0,T]$, subject to the constraint $X(t)\in\mathsf{SDiff}(D)$, is the solution of the following Newton’s equation:
$\rho_{0}\ddot{X}=-\nabla p(t,X),\quad u=\dot{X}\circ X^{-1},\quad p=-\Delta_{X_{\sharp}\rho_{0}}^{-1}\nabla\cdot(u\cdot\nabla u).$ (21)
Which is the Lagrangian version of (1).
Proof. Thanks to the Lemma 2.1, differentiating constraint $X(t)\in\mathsf{SDiff}(D)$ twice leads to the following second-order condition:
$\mathsf{Tr}[(\nabla X)^{-1}(t,a)\nabla\dot{X}(t,a)]^{2}+\mathsf{Tr}[(\nabla X)^{-1}(t,a)\nabla\ddot{X}(t,a)=0$ (22)
Introducing the Lagrange multiplier $\chi(t,a)$ of the above condition, we define
$U(X):=\sup_{\lambda}\int_{D}\chi(t,a)\big[-\mathsf{Tr}[(\nabla X)^{-1}(t,a)\nabla\dot{X}(t,a)]^{2}+\mathsf{Tr}[(\nabla X)^{-1}(t,a)\nabla\ddot{X}(t,a)]\big]da$
Clearly $U(X)=0$ when $X\in\mathsf{SDiff}(D)$ and $U(X)=\infty$ if $X\notin\mathsf{SDiff}(D)$. Now consider the following augmented Appellian functional:
$\tilde{\mathcal{A}}(\ddot{X}(t),\dot{X}(t),X(t))=\int_{D}\frac{1}{2}\lvert\ddot{X}(t,a)\rvert^{2}\rho_{0}(a)+U(X(t,a))da$ (23)
and associated saddle-point problem:
$\inf_{X\in\mathsf{SDiff}(D)}\int_{D}\frac{1}{2}\lvert\ddot{X}(t,a)\rvert^{2}\rho_{0}(a)da=\inf_{X\in\mathsf{SDiff}(D)}\int_{D}\frac{1}{2}\lvert\ddot{X}(t,a)\rvert^{2}\rho_{0}(a)+U(X(t,a))da.$
$\inf_{X\in\mathsf{SDiff}(D)}\sup_{\chi}\int_{D}\frac{1}{2}\lvert\ddot{X}\rvert^{2}\rho_{0}(a)+\chi\big(-\mathsf{Tr}[(\nabla X)^{-1}\nabla\dot{X}]^{2}+\mathsf{Tr}[(\nabla X)^{-1}\nabla\ddot{X}]\big)da.$
The above function is quadratic in $\ddot{X}$ and linear in $\chi$, since we notice that the linear term in $\ddot{X}$ can be rewritten as
$\int_{D}\chi(t,a)\mathsf{Tr}[(\nabla X)^{-1}\nabla\ddot{X}](a)da$
$\int_{D}\chi(t,a)\mathsf{Tr}[\nabla(\ddot{X}\circ A)(X(t,a))]da$
$\int_{D}\chi(t,A(a))\nabla\cdot(\ddot{X}\circ A)(a)da$
$\int_{D}\nabla(\chi\circ A)(X(a))\ddot{X}(a)da.$
Thus, the original saddle-point problem can be reformulated as follows:
$\inf_{X\in\mathsf{SDiff}(D)}\sup_{\chi}\int_{D}\frac{1}{2}\lvert\ddot{X}\rvert^{2}\rho_{0}(a)+\nabla(\chi\circ A)(X(a))\ddot{X}(a)+\chi\mathsf{Tr}[(\nabla X)^{-1}\nabla\dot{X}]^{2}da.$ (24)
Now, since (24) is quadratic in $\ddot{X}$ and linear in $\chi$, we swap the order of inf and sup and consider
$\sup_{\chi}\inf_{X\in\mathsf{SDiff}(D)}\int_{D}\frac{1}{2}\lvert\ddot{X}\rvert^{2}\rho_{0}(a)+\nabla(\chi\circ A)(X(a))\ddot{X}(a)+\chi\mathsf{Tr}[(\nabla X)^{-1}\nabla\dot{X}]^{2}da.$
Thanks to the quadratic formula, maximizing in $X$ yields
$\ddot{X}(t,a)=-\frac{\nabla(\chi\circ A)(X(a))}{\rho_{0}}$ (25)
Now we aim to find
$\sup_{\chi}\int_{D}-\frac{1}{2\rho}\lvert\nabla(\chi\circ A)\rvert^{2}+(\chi\circ A)\mathsf{Tr}[(\nabla X)^{-1}\nabla\dot{X}]^{2}(A(a))da$ (26)
Denote by $\tilde{\chi}=\chi\circ A$, simple algebraic manipulation yields
$\int_{D}\tilde{\chi}\mathsf{Tr}[(\nabla X)^{-1}\nabla\dot{X}]^{2}(A(a))da=\int_{D}\mathsf{Tr}[(\nabla u)^{2}(a)]\tilde{\chi}(a)da=\int_{D}\nabla\cdot(u\cdot\nabla u)(a)\tilde{\chi}(a)da.$
and
$\int_{D}-\frac{1}{2\rho}\lvert\nabla(\chi\circ A)\rvert^{2}da=-\frac{1}{2}\int_{D}\rho^{-1}\nabla\tilde{\chi}\cdot\nabla\tilde{\chi}da=\frac{1}{2}\int_{D}\Delta_{\rho}\tilde{\chi}\cdot\tilde{\chi}da.$
Hence, (26) becomes a quadratic problem in $\tilde{\chi}$. The maximizer is computed similarly as follows:
$\Delta_{\rho}\tilde{\chi}=-\nabla\cdot(u\cdot\nabla u)\quad\Longrightarrow\quad\tilde{\chi}=-\Delta_{\rho}^{-1}\nabla\cdot(u\cdot\nabla u)=p.$ (27)
Hence, we combine (25) and (27) and conclude the desired result. ∎
@@ 3 Hamiltonian and Vorticity Formulation
In this section, we analyze how the above geometric picture flips to the Hamiltonian side. Recall that the Hamiltonian formulation of the motion of ideal fluids reads coadjoint motion of one-form on $\mathfrak{X}_{\sigma}^{*}(D)$, lies in the framework of general Lie-Poisson equations [MR99], while the Leray projector essentially becomes the musical isomorphism from cotangent to tangent vectors. Identifying $\mathfrak{X}_{\sigma}(D)$ with the space of scalar(vector, resp.) potentials by $\nabla^{\perp}(\nabla\times$,resp.) in $2(3$-d,resp.), we obtain the vortex dynamics. Here we extend the above geometric picture to the inhomogeneous setting. A Hamilton-Pontryagin type variational principle and a companion Lagrangian-Eulerian formulation is established. We then establish the vorticity formulation and appreciate such formulation to prove local existence of strong solutions.
@@ 3.1 Hamilton-Pontryagin Principle and Lagrangian-Eulerian Formulation
We now lift ourselves to the cotangent bundle of $\mathsf{SDiff}(D)$ where the Hamiltonian equation naturally lives and we’ll derive it in our setting. To this end, we first write down the kinetic energy as a functional on tangent bundle:
$L:(X,\dot{X})\in T_{X}\mathsf{SDiff}(D)\to L(X,\dot{X}):=\frac{1}{2}\int_{D}\lvert\dot{X}(t,x)\rvert^{2}\rho_{0}(x)dx$
It’s not hard to show that the action-minimizing trajectory augmented by the incompressibility constraint $\det(\nabla X)=1$ is precisely the equation (1). We refer the interested readers to [[LLP11]] for study of relaxed geodesic problem in the sense of Brenier’s generalized flow (cf. [Br89]).
@@
Theorem 3.1(Hamilton-Pontryagin Principle)
Fix $T>0$, assume triplet $(X,u,\xi)\in C^{1}([0,T];\mathsf{SDiff}(D)\times\mathfrak{X}_{\sigma}(D)\times\mathfrak{X}(D))$ is a critical point of the augmented action
$\Phi(u,X,\xi)=\int_{0}^{T}L(u_{t},X_{t})-\langle\xi_{t},\dot{X}-u(X)\rangle_{L_{\rho_{0}}^{2}}dt$ (29)
@@
    Consider the following Hamiltonian:
@@
    Then $(X,u,\xi)$ solves the associated Hamiltonian equation
@@
    Given initial data $(u_{0},\xi_{0})$ satisfying compatibility condition, the solution of (30) is given by the following Lagrangian-Eulerian system:
    $\left\{\begin{aligned} &\dot{X}=u(t,X),A=X^{-1}\\ &\rho(t,\cdot)=\rho_{0}\circ A_{t}\\ &u=\mathbf{P}_{\rho}\bigg[\nabla^{*}A_{t}\bigg(\xi_{0}-\frac{\nabla\log\rho_{0}}{2}\int_{0}^{t}\lvert\dot{X}_{\tau}(x)\rvert^{2}d\tau\bigg)\circ A_{t}\bigg]\end{aligned}\right$. (31)
Proof. (I) As a reminder of our notation, we denote by $A=X^{-1}$ and $\rho:=\rho_{0}\circ A$. We compute the first variation of $\Phi$. We consider smooth perturbations $\delta u\in C^{1}([0,T];\mathfrak{X}_{\sigma}(D))$, $\delta X\in C_{c}^{1}([0,T];\mathsf{SDiff}(D))$ and $\delta\xi\in C^{1}([0,T];\mathfrak{X}_{\sigma}^{*}(D))$ of $\Phi$, respectively, and compute the corresponding variational equations as follows:
@@
    $\frac{d}{d\varepsilon}\bigg\rvert_{\varepsilon=0}\Phi(u+\varepsilon\delta u,X,\xi)=\int_{0}^{T}\int_{\mathbb{R}^{d}}\langle u\rho,\delta u\rangle-\langle\xi,\delta u(X)\rangle_{L_{\rho_{0}}^{2}}dxdt.$ (32)
    Hence vanishing of (32) implies that $\rho(u-\xi\circ A)$ is a gradient, i.e. there exists some scalar function $q$ such that
    $u\rho=(\rho_{0}\xi)\circ A+\nabla q$ (33)
@@
    $\int_{0}^{T}\int_{D}\frac{d}{d\varepsilon}\bigg\rvert_{\varepsilon=0}\frac{1}{2}\lvert u\rvert^{2}\rho_{0}\circ X_{\varepsilon}^{-1}+\langle\xi,\dot{X}+\varepsilon\dot{\delta X}\rangle_{L_{\rho_{0}}^{2}}-\langle\xi,u(X+\varepsilon\delta X)\rangle_{L_{\rho_{0}}^{2}}dxdt$ (34)
@@
    $0=\frac{d}{d\varepsilon}(X_{\varepsilon}\circ X_{\varepsilon}^{-1})=\delta X\circ X_{\varepsilon}^{-1}+\nabla X_{\varepsilon}\circ X_{\varepsilon}^{-1}\cdot\frac{d}{d\varepsilon}X_{\varepsilon}^{-1}$
    hence we conclude:
    $\frac{d}{d\varepsilon}\bigg\rvert_{\varepsilon=0}X_{\varepsilon}^{-1}=-[(\nabla X_{\varepsilon})^{-1}\delta X]\circ X_{\varepsilon}^{-1}\bigg\rvert_{\varepsilon=0}=-[(\nabla X)^{-1}\delta X]\circ A$ (35)
@@
    $\int_{0}^{T}\int_{\mathbb{R}^{d}}\langle\rho_{0}\dot{\xi}+\rho_{0}\nabla^{*}u(X)\xi+\frac{1}{2}\lvert u\circ X\rvert^{2}\nabla\rho\circ X,\delta X\rangle dxdt.$ (36)
    Hence, sending (36), the variational equation in $X$ reads
    $\rho_{0}[\dot{\xi}+\nabla^{*}u(X)\xi]+\frac{1}{2}\lvert u\circ X\rvert^{2}\nabla\rho\circ X=0$ (37)
@@
    $\left\{\begin{aligned} &\dot{X}=u(X),\quad A=X^{-1}\\ &\rho_{0}[\dot{\xi}+\nabla^{*}u(X)\xi]+\frac{1}{2}\lvert u\circ X\rvert^{2}\nabla\rho\circ X=0\\ &u\rho=(\rho_{0}\xi)\circ A+\nabla q,\quad\nabla\cdot u=0\\ &\rho=X_{*}\rho_{0}=\rho_{0}\circ A\end{aligned}\right$. (39)
@@
$\partial_{t}(u\rho-\nabla q)+u\cdot\nabla(u\rho-\nabla q)+\nabla^{*}u\cdot(u\rho-\nabla q)+\frac{1}{2}\lvert u\rvert^{2}\nabla\rho=0$

### v3

@@ 2.3 Geodesic Equation
We prove in this section the following theorem:
Theorem 2.4.
Let $X$ be a geodesic on infinite dimensional Riemannian manifold $(\mathsf{SDiff}(D),\mathsf{G})$ with $X(0)=\mathbf{Id}$. Then, the following statements hold:
  • (I)
    There exists a pressure field $p$ such that
    $\rho_{0}\ddot{X}=-\nabla p\circ X.$ (12)
  • (II)
    Let $u$ be the velocity field of $X$, then
    $p=-\Delta_{\rho}^{-1}\nabla\cdot(u\cdot\nabla u),\quad\rho=X_{\sharp}\rho_{0}.$ (13)
  • (III)
    For $u=\dot{X}\circ X^{-1}$ and $\rho=X_{\sharp}\rho_{0},$ $(u,\rho,p)$ solves the IIE (1).
Proof of Theorem 2.4.
  • (I)
    By D’Alembert principle, we have $\ddot{X}\in T_{X}^{\perp}\mathsf{SDiff}(D)$. Hence, there exists a scalar function $\lambda:[0,T]\times D\to\mathbb{R}$ such that for any $x\in D$:
@@
    $(\nabla^{*}X)^{-1}(x)\int_{D}\nabla^{*}X(z)\nabla p(X(z))\delta_{x}(z)dz=-\nabla p(t,X),$
  • (II)
@@
    Differentiating (15) in time and denoting by $u=\dot{X}_{t}\circ X_{t}^{-1}$, we have for any fixed $a\in D$:
@@
  • (III)
    The above geodesic equation (12) in Eulerian coordinates reads:
@@ 2.4 Second Fundamental Form and Curvature Tensor
Proposition 2.5.
Let $\rho\in\mathsf{Dens}(D)$ and smooth vector field $w\in\mathfrak{X}(D)$, then the following statements hold:
  • (I)
    There exists a unique div-free vector field $v$ and $p\in C^{\infty}(D)$ such that
    $w=\rho v+\nabla p.$ (18)
    moreover, $v=(\mathbf{Id}-\mathbf{Q}_{\rho})(\rho^{-1}w)=:\mathbf{P}_{\rho}(\rho^{-1}w)$. Here, $v$ is not necessarily mean-zero, but $\rho v\in\mathfrak{X}_{0}(D)$ iff $w\in\mathfrak{X}_{0}(D)$.
  • (II)
    Let $\mathcal{H}_{\rho}:=\{\rho^{-1}\nabla\varphi:\varphi\in C^{\infty}(D)\}$, then we have $\mathsf{Ker}(\mathbf{Id}-\mathbf{Q}_{\rho})=\mathcal{H}_{\rho}$. In particular, we have $\mathcal{H}_{\rho}\perp\mathsf{Ran}(\mathbf{P}_{\rho})$ in the sense of $L_{\rho}^{2}$.
Proof of Proposition 2.5.
  • (I)
    Given vector field $w$, for any given $\rho\in\mathsf{Dens}(D)$ we write
    $w=\rho\underbrace{[\rho^{-1}w-\rho^{-1}\nabla\Delta_{\rho}^{-1}\nabla\cdot(\rho^{-1}w)]}_{:={\mathbf{P}}_{\rho}(\rho^{-1}w).}+\underbrace{\nabla\Delta_{\rho}^{-1}\nabla\cdot(\rho^{-1}w)}_{=\rho\mathbf{Q}_{\rho}(\rho^{-1}w)}$ (19)
    Let $v=\mathbf{P}_{\rho}(\rho^{-1}w)$ and $p=\Delta_{\rho}^{-1}\nabla\cdot(\rho^{-1}w)$, we find the decomposition. Moreover, assume $w=\rho v^{\prime}+\nabla p^{\prime}$ for some div-free $v^{\prime}$ and $p^{\prime}\in C^{\infty}(D)$, then $p^{\prime}$ can be recovered from solving the following elliptic problem:
@@
    Uniform ellipticity of $\Delta_{\rho}$ implies the uniqueness of solution $p^{\prime}$, hence the uniqueness of $v^{\prime}$.
  • (II)
    Picking any $\varphi\in C^{\infty}(D)$, we compute $\mathbf{P}_{\rho}(\rho^{-1}\nabla\varphi)$ as follows:
    $\mathbf{P}_{\rho}(\rho^{-1}\nabla\varphi)=\rho^{-1}\nabla\varphi-\rho^{-1}\nabla\Delta_{\rho}^{-1}\nabla\cdot(\rho^{-1}\nabla\varphi)=0,$
    which guaranties $\mathcal{H}_{\rho}\subset\mathsf{Ker}(\mathbf{P}_{\rho})$. Conversely, assume we have $\mathbf{P}_{\rho}\xi=0$, then by the decomposition:
    $\xi=\mathbf{Q}_{\rho}\xi=\rho^{-1}\nabla\Delta_{\rho}^{-1}\nabla\cdot\xi\in\mathcal{H}_{\rho}.$
    which completes the proof of (II).
@@
Theorem 2.6.
The second fundamental form $\Pi_{X}$ of $\mathsf{SDiff}(D)$ at $X$ is given by
@@
Proof of Theorem 2.6.
Formula (20) follows from straightforward computation. Mimicking the definition for second fundamental form in finite dimension:
@@
Remark 2.7.
$\Pi_{X}$ is symmetric on $T_{X}\mathsf{SDiff}(D)\times T_{X}\mathsf{SDiff}(D)$, since
@@
Here, we use the fact that $\nabla\cdot[u,v]=0$ provided $u,v\in\mathfrak{X}_{\sigma}(D)$.
@@
$\mathcal{C}_{\xi,\eta}=\int_{D}[\Pi_{X}(\xi,\xi)(z)\cdot\Pi_{X}(\eta,\eta)(z)-\lvert\Pi_{X}(\xi,\eta)(z)\rvert^{2}]\rho_{0}(z)dz$
@@ 3 Hamiltonian and Vorticity Formulation
In this section, we flip to the Hamiltonian side of the above geometric picture of IIE and derive the Lagrangian-Eulerian formulation of IIE from Hamilton’s action principle. The Hamiltonian formulation of the motion of ideal fluids reads coadjoint motion of one-form on $\mathfrak{X}_{\sigma}^{*}(D)$, lies in the framework of general Lie-Poisson equations [30], while the Leray projector essentially becomes the musical isomorphism from cotangent to tangent vectors (see [31][1]). Identifying $\mathfrak{X}_{\sigma}(D)$ with the space of scalar(vector, resp.) potentials by $\nabla^{\perp}(\nabla\times$,resp.) in $2(3$-d,resp.), we obtain the vortex dynamics.
Here, we extend the above geometric picture to the inhomogeneous setting. A Hamilton-Pontryagin type variational principle and a companion Lagrangian-Eulerian formulation is established, which generalizes the Weber’s formula for perfect fluids in [12]. We then establish the vorticity formulation for IIE in simply connected domain and torus respectively. The reason why we need to treat torus case separately is because the mean value of $\rho u$ is conserved, instead of $u$, hence a global stream function representation for $u$ is not applicable in torus case. A proof of local existence of strong solutions for IIE is obtained based on our vorticity formulation.
@@ 3.1 Hamilton-Pontryagin Principle and Lagrangian-Eulerian Formulation
We now lift ourselves to the cotangent bundle of $\mathsf{SDiff}(D)$ where the Hamiltonian equation naturally lives and we will derive it in our setting. To this end, we first write down the kinetic energy as a functional on tangent bundle:
$L:(X,\dot{X})\in T_{X}\mathsf{SDiff}(D)\to L(X,\dot{X}):=\frac{1}{2}\int_{D}\lvert\dot{X}(t,x)\rvert^{2}\rho_{0}(x)dx$ (21)
It is not hard to show that the action-minimizing problem of $L$, augmented by the incompressibility constraint $\det(\nabla X)=1$, is precisely solved by the solution of equation (1). We refer the interested readers to [29] for study of a relaxed geodesic problem in the sense of Brenier’s generalized flow (cf. [7], [8]).
@@
Theorem 3.1 ((Hamilton-Pontryagin Principle)).
Fix $T>0$ and $\rho_{0}\in\mathsf{Dens}(D)$ satisfying assumption 2.1, we assume the triplet
$(X,u,\xi)\in C^{1}([0,T];\mathsf{SDiff}(D)\times\mathfrak{X}_{\sigma}(D)\times\mathfrak{X}(D))$
is a critical point of the augmented action
$\Phi(u,X,\xi)=\int_{0}^{T}L(u,X)-\langle\xi,\dot{X}-u(X)\rangle_{L_{\rho_{0}}^{2}}dt.$ (23)
@@
    Consider the following controlled Hamiltonian:
@@
    Then $(X,u,\xi)$ solves the associated Hamiltonian ODE:
@@
    Given the initial data $(u_{0},\xi_{0})$ satisfying compatibility condition
    $u_{0}=\mathbf{P}_{\rho_{0}}\xi_{0},\quad\text{equivalently }\xi_{0}=u_{0}+\rho_{0}^{-1}\nabla q_{0}.$
    the solution of (24) is given by the following Lagrangian-Eulerian system:
    $\left\{\begin{aligned} &\dot{X}_{t}=u(t,X_{t}),A_{t}=X_{t}^{-1};\\ &\rho(t,\cdot)=\rho_{0}\circ A_{t};\\ &u_{t}=\mathbf{P}_{\rho}\bigg[\nabla^{*}A_{t}\bigg(\xi_{0}-\frac{\nabla\log\rho_{0}}{2}\int_{0}^{t}\lvert\dot{X}_{\tau}(x)\rvert^{2}d\tau\bigg)\circ A_{t}\bigg].\end{aligned}\right$. (25)
Proof of Theorem 3.1.
  • (I)
    As a reminder of our notation, we denote by $A=X^{-1}$ and $\rho:=\rho_{0}\circ A$. We compute the first variation of $\Phi$. We consider smooth perturbations $\delta u\in C^{1}([0,T];\mathfrak{X}_{\sigma}(D))$, $\delta X\in C_{c}^{1}([0,T];\mathsf{SDiff}(D))$ and $\delta\xi\in C^{1}([0,T];\mathfrak{X}(D))$ of $\Phi$, respectively, and compute the corresponding variational equations as follows:
@@
      $\frac{d}{d\varepsilon}\bigg\rvert_{\varepsilon=0}\Phi(u+\varepsilon\delta u,X,\xi)=\int_{0}^{T}\int_{D}\langle u\rho,\delta u\rangle-\langle\xi,\delta u(X)\rangle_{L_{\rho_{0}}^{2}}dxdt.$ (26)
      Hence, vanishing of (26) implies that $\rho(u-\xi\circ A)$ is a gradient, i.e. there exists some scalar function $q$ such that
      $u\rho=(\rho_{0}\xi)\circ A+\nabla q.$ (27)
@@
      $\int_{0}^{T}\int_{D}\frac{d}{d\varepsilon}\bigg\rvert_{\varepsilon=0}\frac{1}{2}\lvert u\rvert^{2}\rho_{0}\circ X_{\varepsilon}^{-1}+\langle\xi,\dot{X}+\varepsilon\dot{\delta X}\rangle_{L_{\rho_{0}}^{2}}-\langle\xi,u(X+\varepsilon\delta X)\rangle_{L_{\rho_{0}}^{2}}dxdt.$ (28)
@@
      $0=\frac{d}{d\varepsilon}(X_{\varepsilon}\circ X_{\varepsilon}^{-1})=\delta X\circ X_{\varepsilon}^{-1}+\nabla X_{\varepsilon}\circ X_{\varepsilon}^{-1}\cdot\frac{d}{d\varepsilon}X_{\varepsilon}^{-1}.$
      Hence, we have:
      $\frac{d}{d\varepsilon}\bigg\rvert_{\varepsilon=0}X_{\varepsilon}^{-1}=-[(\nabla X_{\varepsilon})^{-1}\delta X]\circ X_{\varepsilon}^{-1}\bigg\rvert_{\varepsilon=0}=-[(\nabla X)^{-1}\delta X]\circ A.$ (29)
@@
      $\int_{0}^{T}\int_{D}\langle\rho_{0}\dot{\xi}+\rho_{0}\nabla^{*}u(X)\xi+\frac{1}{2}\lvert u\circ X\rvert^{2}\nabla\rho\circ X,\delta X\rangle dxdt.$ (30)
      By arbitrariness of $\delta X$ in (30), the variational equation in $X$ reads
      $\rho_{0}[\dot{\xi}+\nabla^{*}u(X)\xi]+\frac{1}{2}\lvert u\circ X\rvert^{2}\nabla\rho\circ X=0.$ (31)
@@
      $\left\{\begin{aligned} &\dot{X}=u(X),\quad A=X^{-1};\\ &\rho_{0}[\dot{\xi}+\nabla^{*}u(X)\xi]+\frac{1}{2}\lvert u\circ X\rvert^{2}\nabla\rho\circ X=0;\\ &u\rho=(\rho_{0}\xi)\circ A+\nabla q,\quad\nabla\cdot u=0;\\ &\rho=X_{\sharp}\rho_{0}=\rho_{0}\circ A.\end{aligned}\right$. (33)
@@
    $\partial_{t}(u\rho-\nabla q)+u\cdot\nabla(u\rho-\nabla q)+\nabla^{*}u\cdot(u\rho-\nabla q)+\frac{1}{2}\lvert u\rvert^{2}\nabla\rho=0.$